UT Sep 25 2022 - Night #14 of Restart - Split night with PETS-PEPSI




Night Info

  • Sunset 18:13
  • Nautical Twilight Ends 19:06
  • Astro Twilight Ends 19:35
  • Astro Twilight Begins 04:48
  • Nautical Twilight Begins 05:17
  • Sunrise 06:10
  • Moonset 18:02
  • Moonrise 05:45
  • Moon illumination ~0%
  • PETS/PEPSI - at start of night
  • Handover to LBTO ~01:27

Staff

  • OSA: D.G.Huerta
  • ISp: Mtn Staff
  • Mtn Manager: R. Hannsen
  • ISA/SO: A.Becker
  • AO: D. Miller (on-call)

Night Plan

ALTA prediction for Tonight

20220924_see_evol_time_Height_20_20000.png

Starting Config: PEPSI-PFU Bino

Priorities are
  • Basic checkout of MODS
  • AGW1/2 Transforms
  • Field Aberration

MODS On-Sky Checkout

AGW 5/6 Checkout
  • Acquire on pointing star
  • Collimation check
  • Measure position of WFS hotspot in twilight
    • Check SH illumination
      • made backup copies of /home/telescope/TCS/Configuration/GCS/MODS_L.cfg andMODS_R.cfg
      • changed maxOffsetToHotspot and maxHotspotFromProbe to 2.54
      • Use the arrow controls in the GCSGUI eng panel
  • check MODS1/MODS2 guide probe wrt/rotator center
    • Files needed:
    • agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
    • PLEASE NOTE:
      • Y differential can be as much as 40 pixels in Y -this is a hardware limit - do not adjust Y if ~40 pixels or less off
MODS Verification
  • Scripts: /home/lbto/Service_Observing/Restart_2022/MODS/MODS_Checkout_V2.6.3.xml
    • OT file may be updated before sunset or handover
  • DONE MODS Rotator Center measurement - high (0.5 hr)
    • Requires pointing mask (facility mask)
    • Take pairs of images 180 degrees apart
    • Pick visible target - min is 1 pair 180 deg apart
    • Scripts generated for M52 (end-of-night) and M71 (beginning). Use these, since I have hand-edited the slitmask to be the PointingMask.
  • DONE MODS AGW 5/6 - Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
    • Target Masks:
      • ID 532732 - mods.4.g04d25ep.mms
      • ID 531713 - mods.2.g00d25ep.mms
      • ID 531956 - mods.3.g2025ep2.mms
      • MMS files have been uploaded to /home/lbto/Service_Observing/Restart_2022/MODS/mms
    • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1
      • The files needed are
      • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Rotational offsets needed to align MODS MOS masks (2021B to 2022A):
instr mean median sig_n-1 n
MODS1 -0.03 -0.03 0.05 33
MODS2 -0.05 -0.04 0.05 55
dtheta_2021_to_2022.png
  • MODS AGW Transform data -Medium (High only if RotZP change) 1hrs
    • Not needed
    • If RotZP is > 0.1 deg then proceed with transform otherwise skip
      • if time, take the transform (after other tasks completed) - 12 points
    • must do RotZP First
    • Seeing 1.2" or better
    • Targets
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
  • Spectrophotometric standard all modes -High (to compare SX vs DX) - Critical - 1hr-2hr
    • If non-photometric - take 1 image and go through as many modes in order of priority below, if photometric - 3 images
    • Pick any spectro-photometic standard star visible
      • Dual Grating
      • Red Grating
      • Blue Grating
      • Dual Prism
      • Red Prism
      • Blue Prism
    • Scripts for G191-B2B and GD71 are in the directory, for both Nimgs=3 and Nimgs=1. G191-B2B will take about 45-min (Nimgs=3) or 15-min (Nimgs=1) not counting acquisition.
  • Test ScienceOB - high 1hr
    • If low on-time move to AGW1/2 Transform
    • If seeing conditions too poor for AGW Transform take test science OB
    • Any conditions
      • LS
        • Arp 230 (dual grating simple bright target - any conditions observable) - visible 22:30-04hr
          • poor weather science script already prepped w/finder chart
      • Imaging
        • Target?
        • Could be photometric fields if clear
          • sdss_r0329_d39 field is one with a good number of SDSS DR12 stars, also sdss_r0340_d40
        • Could be NS target
Other MODS Checks
  • Lower priority relative to other MODS tests and taking AGW Transform confirimatory data and Field Aberration Confirmatory Data
    • Photometric Standards
      • any seeing, but must be clear
        • sdss_r0329_d39 field is one with a good number of SDSS DR12 stars, also sdss_r0340_d40
    • Skyflats - Medium

LUCI AGW1/2 Transform Confirmation

  • Collect transform data - (2hrs) - confirmatory - Requires LUCIs - transforms done and will be installed before sunset
    • update oaccontrol.conf with new results from D.Miller
      • update oaccontrol.conf with new results from D.Miller
      • backup file first
      • add the snippets from D.Miller in the correct place for AGW1 and AGW2
      • rdwrconfig -u1 and rdwrconfig -u2
      • need to restart PCS and GCS to take hold
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
      • 04 30 27 -00 36 22 Stone_D_LUCI_PA0_short.list (12 targets)
    • Note the LUCI frame numbers associated with each offset in IDL script

Field Aberration Confirmatory Data

  • Reconfigure for bino-LUCIs
  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour) - update installed already
  • Can be done in parallel (data collection, analysis in serial)
  • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
  • 1.2" or better seeing (1" or better ideal)
  • Have OSA go to field closeby, point and collimate, then preset to target
  • Target Fields
    • BS9177 18:39:33.7 +49:05:38.0
    • StoneP1 22:37:45.0 -01:13:40.45
    • BS9188 23:30:33.4 +38:18:57.0
    • BS9107 03:32:03.0 +37:20:40.0

Backup Poor Weather

TMS Testing/LBCs
  • Target? (dense star field, e.g. one at low Galactic latitude, to assess IQ, or another scientifically valuable field)
    • Any OB in the LBC restart script is fine as long as there are lots of stars --- OB added for gal+ecl field.
    • ONLY GOAL - make sure TMS works without any issues - IQ is of seconary importance
LUCI
  • LUCI Tasks - once poor weather MODS are done
    • LUCI Scripts
      • LUCI_2022_Restart_V1.81.xml - all scripts (except LS Rot ZP)
      • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards (fixes AGW patrol problem)
    • Photometric standards
      • must be clear, any seeing
    • Imaging check (1hr) - medium priority
      • ANY seeing OK
      • dither in-field
        • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am - only need to run 4 dithers
        • dither out-of-field (run one)
      • dither out-of-field (run one)
        • M52 - JHK, dither 5' out of field - visible all night
        • M40 - Broad filters, few narrow filters - visible all night

Night Log

Googley Doc Log Link

-- BarryRothberg - 23 Sep 2022
I Attachment Action Size Date Who Comment
20220924_see_evol_time_Height_20_20000.pngpng 20220924_see_evol_time_Height_20_20000.png manage 121 K 24 Sep 2022 - 21:02 BarryRothberg  
dtheta_2021_to_2022.pngpng dtheta_2021_to_2022.png manage 15 K 25 Sep 2022 - 08:05 OlgaKuhn Rotational offsets needed to align MODS MOS masks (2021B to 2022A)
galplane.pngpng galplane.png manage 129 K 23 Sep 2022 - 23:52 OlgaKuhn Equatorial coordinates of Galactic plane (for choosing dense star fields for LBC+TMS checks)
Topic revision: r8 - 25 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback